(Bond valuation)  ​Fingen's 14​-year, $1,000 par value bonds pay 8 percent interest annually. The market price of the bonds is ​$1,130 and the​ market's required yield to maturity on a​ comparable-risk bond is 5 percent.

a.  Compute the​ bond's yield to maturity.

b.  Determine the value of the bond to​ you, given your required rate of return.

c.  Should you purchase the​ bond?

round to two decimal places.
Answer #1

a.  Compute the​ bond's yield to maturity.

Yield To Maturity(YTM) = (interest per annum + (Redemption price - Current market price) / life remaining to maturity) / (.4*Redemption price)+ (.6*Current market price)

= (1000*8%)+)(1000-1130)/14)/(.4*1000+.6*1130)

= (80-9.28571428571)/1078

= 6.55%

b.  Determine the value of the bond to​ you, given your required rate of return.

Bond Valuation: The value of bond is the present value of the expected cashflows from the bond,discounted at Yield to Maturity(YTM).

Year Cash flow PVAF/PVF@5% Present Value (Cashflow*PVAF/PVF)
1-14 80 9.8986 791.89
14 1000 0.5051 505.07

Current Market Price of Bonds = Cashflow*PVAF/PVF

= 791.89+505.07

= $1296.96

*PVAF = (1-(1+r)^-n)/r

**PVF = 1 / (1+r)^n

c.  Should you purchase the​ bond?

Here actual price ($1,130) < Fair price ($1,296.96), so the bond is undervalued, purchase the bonds.